find the mass of magnesium oxide that can be formed by burning completely 30g of magnesium ribbon
1
Expert's answer
2020-05-19T08:58:05-0400

2Mg+O2 => 2MgO

n(Mg)=m(Mg)/Mr(Mg)=30/24=1,25 moll

n(Mg)=n(MgO)

=>m(MgO)=n(MgO)*Mr(MgO)=1,25*40=50g